OpenStudy (mopsy510):

The radius of a circle is 20 miles. What is the angle measure of an arc bounding a sector with area 10π square miles?

OpenStudy (anonymous):

Area of a sector = n/360 (pi)(radius)^2

OpenStudy (ranga):

Knowing the radius find the area of the circle. A full circle is 360 degrees. So 360 degrees is equivalent to full area of the circle x degree is equivalent to 10(pi) It is a proportion problem. Solve for x.

OpenStudy (mopsy510):

So 360-10(pi)?

OpenStudy (ranga):

360 / ((pi)(r^2)) = x/(10pi)

OpenStudy (mopsy510):

Is the answer \[\approx 1\]

OpenStudy (mopsy510):

Not 1 i meant 11.4

OpenStudy (anonymous):

mopsy...?

OpenStudy (ranga):

x / 10pi = 360 /((pi)(r^2)) To isolate x, multiply both sides by 10pi x = 360 /((pi)(r^2)) (10pi) = 3600/r^2 = 3600 / 20^2 = 9 degrees